OpenStudy (anonymous):

Rewrite the following linear equation in slope-intercept form. Write your answer with no spaces. y+4 = -2(x-1)

OpenStudy (anonymous):

y+4 = -2(x-1) First put the y by itself. y=-2(x-1)-4 Then distribute your -2. y=-2x+2-4 Combine like terms after that. y=-2x-2 Then thats it. You cant simplify it further.

OpenStudy (anonymous):

no What form is it in

OpenStudy (anonymous):

rise-run,point slope,slope intercept,standard ?

OpenStudy (anonymous):

The equation you wrote is in point slope form. y=-2x-2 is slope intercept since it is in the form y=mx+b.

OpenStudy (anonymous):

Thank you